OpenStudy (anonymous):

How many solutions exist for the system, x + y = 1 and y = -x + 1?

OpenStudy (anonymous):

One

OpenStudy (anonymous):

ONLY ONE PAIR OF SOLUTION EXISTS @gabbie96

OpenStudy (debbieg):

Hmmmmm..... nope, look again. Notice anything "fishy" about those two equations?

OpenStudy (anonymous):

Ohhh, I read it wrong! They're the same equation, aren't they?

OpenStudy (debbieg):

@gabbie96 , what do you notice?

OpenStudy (anonymous):

you can re-arrange the first equation: x + y = 1 y = 1 - x so we then have y = 1 - x y = -x + 1 it's the same equation.

OpenStudy (debbieg):

One way to see what "kind" of system you have, is to solve both equations for y. If you get the SAME equation, then the system is dependent: it's just one line, so EVERY point on that line is a solution, so infinite solutions.

OpenStudy (debbieg):

If you get two lines with the same slope (but different y-intercepts) then they are parallel lines, so NO solutions, since they never intercept.

OpenStudy (anonymous):

Wouldn't this mean there are an infinite number of solutions? As long as x and y add up to 1 then it works. example: x=-2, y=3 x=1, y=0 I don't know, I could be wrong.

OpenStudy (debbieg):

and if you get two distinct, non-parallel lines (all other cases) then you get one solutions, since there is one point of interception.